Pls help me with this qn: One thousand cubes are arranged to form a large cube. All the surfaces of the large cubes are then painted. What fraction of the cubes are not painted at all?
Answer by checkley71(8403)

WHAT YOU HAVE HERE IS AN INTERESTING PUZZLE.
1000 CUBES FORM A 10x10x10 LARGE CUBE.
THEREFORE THERE ARE 100 CUBES FORMING THE TOP & THE BOTTOM.
THIS ACOUNTS FOR 200 PAINTED CUBES.
NOW FOR THE 4 SIDES WE HAVE ALREADY ACCOUNTED FOR 20 (10 AT THE TOP ROW & ANOTHER 10 ON THE BOTTOM ROW) THAT LEAVES US WITH 80 ON THE 4 SIDES.
HOWEVER THERE ARE 8 EDGE CUBES AS PART OF THE ADJOINING SIDE.
NOW WE HAVE TO SUBTRACT 8 FROM THE 80=72 FOR EACH SIDE BECAUSE OF THE OVERLAPPINGCORNERS.
THUS WE SHOULD NOW HAVE A TOTAL OF 2*100+4*72=488 CUBES THAT HAVE AT LEAST ONE SIDE PAINTED.
THIS LEAVES 1000-488=512 CUBES UNPAINTED.
OR 512/1000=51.2% UNPAINTED CUBES.
WISH I COULD DEVELOPE A SIMPLE FORMULA FOR THIS TYPE OF PROBLEM BUT I'M A PICTORIAL TYPE OF MATHEMATICIAN AND FIND A DIAGRAM MOST EFFECTIVE.
I JUST HAD A BRAIN STORM.
RATHER THAN WORKING WITH THE PAINTED CUBES LET'S LOOK AT THE UNPAINTED CUBES.
THE INNER CORE THAT ISN'T PAINTED MEASURES 8x8x8=512.
IMMAGINE THAT 51.2% ARE CORE CUBES & UNPAINTED.
SHOULD HAVE THOUGHT OF THIS APPROACH FIRST.
OH WELL WHEN YOU'RE WORKING A NEW FIELD YOU DON'T ALWAYS FIND THE EASIEST SOLUTION ON THE FIRST TRY. THIS IS A PRIME EXAMPLE.